OpenStudy (anonymous):

Today is my first day being on this site and I love it, I get help quickly and I really appreciate it. :) I do have one more though... How would I solve -12=3(x+2) ?

OpenStudy (skullpatrol):

Welcome, do you have any ideas how to solve this equation?

OpenStudy (skullpatrol):

$$ -12=3(x+2)$$

OpenStudy (anonymous):

Thank you. I would multiply 3(x) and 3(2) therefore I would get -12=3x + 6 so -12=18x is that right?

OpenStudy (skullpatrol):

3x + 6 does not equal 18x

OpenStudy (anonymous):

I multiplied I should have added. It will be 9x. They are both positive numbers so I would add them. So.. -12=9x

OpenStudy (skullpatrol):

Nope, 3x + 6 does not equal 9x.

OpenStudy (skullpatrol):

3x + 6 is in its simplest form as it is now, it can not be simplified any further, do you know why?

OpenStudy (anonymous):

No I don't. I am reviewing these things and don't remember :(

OpenStudy (anonymous):

I was thinking I would divide both the 3 and the -12 by 3 to get the variable by itself...?

OpenStudy (anonymous):

This is very elementary and now I'm confusing myself further, I do appreciate your help.

OpenStudy (skullpatrol):

Yes, try dividing both sides of the equation by 3 and show me what you get please :)

OpenStudy (skullpatrol):

$$ -12=3(x+2)$$ $$\frac{ -12}{3}=\frac{3(x+2)}{3}$$

OpenStudy (anonymous):

x=-2 ?

OpenStudy (skullpatrol):

What is -12/3?

OpenStudy (skullpatrol):

@nikv00 I would suggest you go to YouTube and look up Khan Academy specifically solving equations and watch them :)

OpenStudy (skullpatrol):

http://www.youtube.com/watch?v=p5e5mf_G3FI

OpenStudy (anonymous):

I will, thanks for your help. :)

OpenStudy (skullpatrol):

np :)